OpenStudy (anonymous):

MEDALLL!!!!! Which number line shows the solution to −3 − (−1)? @ganeshie8

OpenStudy (anonymous):

−3 − (−1) = -3 + 1 = -2

OpenStudy (anonymous):

-3 -(-1) = -3 + 1 = -2

OpenStudy (anonymous):

Is the answer....

OpenStudy (anonymous):